OpenStudy (anonymous):

How can I add these two equations in order to solve for y? y=(1/2x)-1 y=((-2)x)+4

OpenStudy (anonymous):

Multiply each side of the first equation by 4. The add the equations together. The x terms will subtract out.

jhonyy9 (jhonyy9):

how are these exactly ? 1 y = ---x - 1 2 (-2)x y= ----- 4 in this way is right ?

OpenStudy (anonymous):

4 y = (2 x) - 4 y = -2x + 4 y = 0 x = 2